OpenStudy (nuccioreggie):

Convert 7 over 8 to a decimal and tell whether it terminates or repeats. 7.8, a number that terminates 0.875, a number that terminates 7.8 with a repeating bar over 8., a number that repeats 0.875 with a repeating bar over 875., a number that repeats

OpenStudy (ilovecake):

A.

OpenStudy (anonymous):

0.875 is the decimal

OpenStudy (ilovecake):

what? no it is A 7.8, a number that terminates.

OpenStudy (anonymous):

But 7.8 is not the correct decimal. It is 0.875

OpenStudy (ilovecake):

A.

OpenStudy (anonymous):

0.875 Is the correct decimal

OpenStudy (nuccioreggie):

Imm going to check with. @bohotness. To see if thats correct

OpenStudy (anonymous):

It is B

OpenStudy (ilovecake):

A.7.8

OpenStudy (nuccioreggie):

@amistre64. Who is correct?

OpenStudy (amistre64):

what is 7/8? 1/2 = .5 1/4 = .25 1/8 = .125 any multiple of 1/2 or power of it will be terminating

OpenStudy (anonymous):

I don't get it..

OpenStudy (amistre64):

if something repeats, then it goes on forever ... if somethings ends, it terminates

OpenStudy (amistre64):

since 1/2 is a terminating decimal, any multiple of it is terminating and any power of it is terminating

OpenStudy (amistre64):

what is 7*.125?

OpenStudy (nuccioreggie):

0.875

OpenStudy (nuccioreggie):

So it is A @@amistre64

OpenStudy (nuccioreggie):

@amistre64

OpenStudy (nuccioreggie):

I ment so its B @amistre64